OpenStudy (anonymous):

Please Help: Using complete sentences, explain what might happen if the order of operations was used to solve a literal equation.

OpenStudy (anonymous):

depending on the operations used the answer would come out differently. addition: no change multiplication: no change subtraction: change in sign division: change in value

OpenStudy (anonymous):

Thank you....can you solve this? Using complete sentences, explain how solving a literal equation is similar to or different from simplifying an expression such as 6 - 2(52 + 7) ÷ 4

OpenStudy (anonymous):

The order of operations (PEMDAS) is the only correct way to solve: 6 - 2(52 + 7) ÷ 4 = 6 - 2(59) ÷ 4 = 6 - 118 ÷ 4 = 6 - 29.5 = -23.5